Find simpler statement forms that are logically equivalent to p ⊕ p and (p ⊕ p) ⊕ p.

b) Is (p ⊕ q) ⊕ r ≡ p ⊕ (q ⊕ r)? Justify your answer.

c) Is (p ⊕ q) ∧ r ≡ (p ∧ r) ⊕ (q ∧ r)? Justify your answer.


1
Expert's answer
2021-03-17T12:09:17-0400

a) Let us find the truth table for "p \u2295 p" and "(p \u2295 p) \u2295 p:"


"\begin{array}{|c|c|c|c|c|c|c|} 
\hline
p & p\oplus p & (p\oplus p)\oplus p \\
\hline
0 & 0 & 0 \\
\hline
1 & 0 & 1 \\
\hline

\end{array}"


It follows that "p ⊕ p\equiv 0" and "(p ⊕ p) ⊕ p\equiv p".



b) Let us find the truth table for "(p \u2295 q) \u2295 r" and "p \u2295 (q \u2295 r)":


"\begin{array}{|c|c|c|c|c|c|c|} 
\hline
p & q & r & p\oplus q & (p\oplus q)\oplus r & q\oplus r & p\oplus (q\oplus r)\\
\hline
0 & 0 & 0 & 0 & 0 & 0 & 0\\
\hline
0 & 0 & 1 & 0 & 1 & 1 & 1\\
\hline
0 & 1 & 0 & 1 & 1 & 1 & 1\\
\hline
0 & 1 & 1 & 1 & 0 & 0 & 0\\
\hline
1 & 0 & 0 & 1 & 1 & 0 & 1\\
\hline
1 & 0 & 1 & 1 & 0 & 1 & 0\\
\hline
1 & 1 & 0 & 0 & 0 & 1 & 0\\
\hline
1 & 1 & 1 & 0 & 1 & 0 & 1\\
\hline
\end{array}"


Since the formulas "(p \u2295 q) \u2295 r" and "p \u2295 (q \u2295 r)" have the same truth value in all cases, they are logically equivalent. It follows that "(p \u2295 q) \u2295 r \u2261 p \u2295 (q \u2295 r)."


c) Let us find the truth table for "(p \u2295 q) \u2227 r" and "(p \u2227 r) \u2295 (q \u2227 r)":


"\begin{array}{|c|c|c|c|c|c|c|c|} 
\hline
p & q & r & p\oplus q & (p\oplus q)\land r & p\land r & q\land r &p\land r\oplus q\land r\\
\hline
0 & 0 & 0 & 0 & 0 & 0 & 0 & 0\\
\hline
0 & 0 & 1 & 0 & 0 & 0 & 0 & 0\\
\hline
0 & 1 & 0 & 1 & 0 & 0 & 0 & 0\\
\hline
0 & 1 & 1 & 1 & 1 & 0 & 1 & 1\\
\hline
1 & 0 & 0 & 1 & 0 & 0 & 0 & 0\\
\hline
1 & 0 & 1 & 1 & 1 & 1 & 0 & 1\\
\hline
1 & 1 & 0 & 0 & 0 & 0 & 0 & 0\\
\hline
1 & 1 & 1 & 0 & 0 & 1 & 1 & 0\\
\hline
\end{array}"


Since the formulas "(p \u2295 q) \u2227 r" and "(p \u2227 r) \u2295 (q \u2227 r)" have the same truth value in all cases, they are logically equivalent. It follows that "(p \u2295 q) \u2227 r \u2261 (p \u2227 r) \u2295 (q \u2227 r)."